### v1

@@ IX.1 Introduction
Radiative proton capture on (=3/2-) at solar energies proceeds through direct capture to the ground state of (=2+). This capture occurs predominantly at separations well beyond the range of the strong interaction via transitions from and partial waves, as the amplitude and contributions from higher partial waves are negligible in the energy range of interest [1961NucPh..24...89C]. The importance of this reaction for the determination of the high energy solar neutrino spectrum and the experimental data reviewed by SF I inspired several experiments that, taken together, provided a consistent picture of the energy dependence of the reaction cross section, despite discrepancies in absolute scale. Prior to the review of SF II, most direct measurements were performed using a radioactive target and an intense proton beam. In spite of the great care taken in evaluating the common mode errors (CMEs), a remarkable discrepancy persisted even in the most recent experiments, that was finally handled in SF II by inflating the stated experimental uncertainties. Indirect measurements of the cross section of based on Coulomb dissociation were considered, but not included in the final recommendation of SF II for the zero-energy astrophysical factor , whose uncertainty was dominated by the theoretical contribution. Therefore, our analysis focused on these aspects (discrepancy of experimental data, indirect measurements, theoretical uncertainties), as discussed in the next sections.
@@ IX.2.1 Direct Measurements
The thorough review of published work done in SF II investigated the influence of beam-target overlap, target stoichiometry, beam energy loss, and the backscattering of 8B recoils on CMEs and led to the selection of a homogeneous group of well-documented data sets. Since then, no further information became available on these aspects of direct measurements. The data that formed the basis of the SF II recommendation were those of 2003PhRvC..67f5805B, 2003PhRvL..90b2501B, 1998PhRvL..80..928H, 2001PhRvL..86.3985H, 1999PhLB..462..237H, 2001NuPhA.696..219S, and the BE3 data set of 2003PhRvC..68f5803J. Here we adopt the revised BE3 data presented by 2010PhRvC..81a2801J. In addition to the complete set of radiative capture measurements that were the basis of the SF II recommendation for , we consider the only new experiment reported thereafter of 2022PhLB..82436819B, which used a radioactive ion beam and the recoil mass separator ERNA to detect the 8B recoils. While these new results contribute to the determination of with different, well controlled systematics [2018EPJA...54...92B], the relatively large statistical uncertainty of the measurement limits its impact.
New scattering data have been published. Elastic and inelastic scattering cross sections at relative kinetic energies between 474 keV and 2.74 MeV were measured and wave scattering lengths were inferred by 2019PhRvC..99d5807P. At higher energies between 1.6 and 3.4 MeV, thick target elastic and inelastic scattering excitation function measurements were used to infer the existence of new resonances above 1.8 MeV [2013PhRvC..87e4617M]. These high lying resonances have no direct influence on .
@@ IX.2.2 Indirect Measurements
In SF II, Coulomb dissociation measurements based on the formalism developed by 1986NuPhA.458..188B were not included in the determination of due to disagreements over whether the analyses of the various measurements properly accounted for the contributions of transitions. Despite these reservations, the Coulomb breakup data were found to be in agreement with radiative capture measurements. In order to include these consistent data in formulating our recommendation for , we inflated the common mode errors of the Coulomb breakup measurements to adequately account for the uncertainty in the components assumed in each analysis.
The common mode errors are obtained from a linear sum of those given in the manuscripts and an additional common mode error due to the estimated size of the component in each measurement. This additional error is applicable in one direction only, downward for the 1998EPJA....3..213K and 2006PhRvC..73a5806S analyses of the RIKEN and GSI measurements, respectively, and upward for the 2003PhRvC..68d5802D analysis of the Michigan State University (MSU) measurement, according to whether the component was included or excluded. Though the 2003PhRvC..68d5802D analysis did include a common mode error contribution of 2.5% due to the component (which represented 5% of the total cross section), we have included an additional 2.5% contribution to the MSU experiment. The size of the additional common mode errors for the RIKEN and GSI measurements was based on scaling this 2.5% according to the predicted size of the contribution to the cross section in these measurements relative to that in the MSU measurement, evaluated using first-order perturbation theory and the and factors calculated in a potential model of 8B [1996ZPhyA.356..293B]. Thus the additional single-sided common mode errors added to the GSI, MSU, and RIKEN measurements are -2.1%, 2.5%, and -4.3%, respectively.
@@ IX.3 Theory
The size of the theoretical contribution to the uncertainty in the SF II estimate of was twice that of the experimental contribution. Since then, significant progress has been made. A new potential model calculation was performed by 2019NuPhA.983..175D. The large-scale computational demands of calculations using high-quality nucleon-nucleon (NN) and three-nucleon (3N) interactions beyond =5 had hindered the ab initio approach in the past. Nevertheless progress has been rapid and it has been possible to extend ab initio calculations to , albeit without 3N forces. 2011PhLB..704..379N performed such a calculation using the no-core shell model in the continuum with a nucleon-nucleon interaction derived from chiral EFT at fourth order. The calculation converged with respect to the model space size and reproduces experimental data for the factor despite having virtually no free parameters (though the similarity renormalization group evolution parameter was tuned to reproduce the binding energy of 8B). The final value of =19.4(7) eV b agrees with the SF II recommendation and, as will be shown later, with our analysis. More recently, a new approach was introduced to combine experimental measurements with ab initio predictions, resulting in an “ab initio-informed evaluation” Kravvaris23-PLB that arrived at a value of eV b.
@@
2015PhLB..751..535Z and 2018PhRvC..98c4616Z include such an excited contribution. 2022PhRvC.106a4601H provides a calculation without the contribution below and one with it above that also includes an contribution to the resonance. The two groups’ halo EFT calculations with the contributions differ in several respects that are elaborated in 2022PhRvC.106a4601H. Differences in the final bound state calculation affect the treatment of the short-range interactions and divergences in the EFTs. The interpretation of the asymptotic normalization constants (ANCs) in terms of elastic scattering parameters is affected. However, the momentum dependence of the capture cross section is not impacted by the ANCs, which are fitted to capture data. Thus there is no effect on the cross section. In the incoming channel, 2022PhRvC.106a4601H use a more general form of the -wave short-ranged interaction involving the core based on the low-energy symmetry that is included only at next-to-next-to-leading order (NNLO). In contrast, 2015PhLB..751..535Z and 2018PhRvC..98c4616Z include a short-ranged interaction at LO that could make a difference in the momentum dependence of the cross section. However, as shown in an order-by-order calculation [2022PhRvC.106a4601H], the short-ranged interaction is a sub-leading effect so the two halo EFTs should have similar low-momentum dependence. The third important difference is in the treatment of two-body currents, which affects EFT error estimates. 2015PhLB..751..535Z and 2018PhRvC..98c4616Z include this at next-to-leading order (NLO) whereas 2022PhRvC.106a4601H estimate it to be an N3LO effect. Bayesian estimates of the two-body current [2015PhLB..751..535Z, 2018PhRvC..98c4616Z] are not strongly constrained by capture data, consistent with this being a higher-order contribution.
Another widely used approach for describing the cross sections of low-energy nuclear reactions is R-matrix theory. In R-matrix theory the 7Be(B cross section can be parametrized in terms of a few parameters representing either “real” or “background” poles. The real poles correspond to the observed states and resonances, while the background terms correspond to the “mean-field” effects that are of the non-resonant type. In the present case for example, a background pole in the 1- channel is needed so as to reproduce the non-resonant part of the capture. The parameters of the real pole corresponding to the bound state are determined by the binding energy and the two ANCs, while the 1+ resonance parameters are determined by its energy and partial decay widths (both proton and radiative). The R-matrix analysis of this reaction carried out by 1995NuPhA.588..693B arrived at a value of eV b; however, it was based on data that were excluded in SF II and here, with the exception of those of 1983PhRvC..28.2222F.
@@ IX.4 Fitting Procedures
In SF II, it was noted that the different data sets had similar energy dependences. The discrepancies among the data sets were primarily due to different absolute normalizations. We address the normalization issue by introducing scaling factors associated with the CMEs of measurements that arise from the parts of the systematic uncertainties that are independent of energy or only very weakly energy-dependent. This contrasts with the SF II approach, in which the errors were inflated by the factor . The procedure for handling discrepant data sets adopted here is mathematically rigorous and aligns naturally with the method of estimating theoretical uncertainty discussed below. We prefer this approach to the inflation factor method used in SF II, since the absolute scales of the different data sets can be used as constraints for each other, while there is no information available for a reevaluation of the CMEs. The procedure we used, described below, can be applied to both frequentist and Bayesian analyses.
Both minimization (a frequentist approach) and Bayesian posterior probability distribution function evaluation were performed to fit EFTs to the data; in the R-matrix analysis, only the former was used. The likelihood function, defined as the conditional probability for the data given theoretical parameters , enters both fitting procedures. It is expressed as [1994NIMPA.346..306D]
@@ IX.5 Data Analysis and Determination
Bayesian fits were performed with halo EFT expressions from both 2020JPhG...47e4002Z and 2018EPJA...54...89H that are compatible with the various fits. The Bayesian analysis had advantages over the fits. The capture proceeds through the and channels. The reaction is known to be peripheral, resulting in a that has a sub-leading dependence on the strong interaction in the initial state [2000PhRvC..62f5803B]. This behavior was shown to persist away from the threshold in halo EFT, in which for capture the strong interaction in the incoming -wave channel only contributes at NNLO.
@@
The Bayesian fits involved 6 theoretical parameters and 11 normalization constants for the data sets. The priors for the normalization constants were as described earlier. The priors for the theoretical parameters are based on the underlying EFT assumptions [2022PhRvC.106a4601H] about the expansion in a small ratio where represents the low momentum of infrared physics and the cutoff of the theory. The and scattering lengths fm and fm [2019PhRvC..99d5807P] are taken as inputs. For capture at LO, the only fit parameter, the effective momentum is assumed to scale as and accordingly we assume a uniform prior for between -100 MeV/ and 1.5 MeV/, where the upper limit is set by the physical constraint that the ANC squares have to be positive. At NLO, the ANC2 is a fit parameter whose prior is taken from a normal distribution determined by experiment. At NNLO, we use the ab initio calculation of [2018PhRvC..98c4616Z] to draw the ANC2 from a normal distribution. A ratio of the -wave scattering lengths in the coupled-channel - that is assumed to be in the EFT is drawn from a normal distribution with a mean of 0 and a standard deviation of 10 to cover a wide range. The capture is dominated by the transition and it requires three parameters: a -wave scattering volume , an effective momentum and a two-body current to regulate divergences. 2022PhRvC.106a4601H determined and from the the narrow resonance energy and width. Here, we kept fixed while drawing from a uniform distribution between -150 and -50 MeV. The Bayesian fit gave . The two-body coupling was drawn from a uniform distribution and the fits gave a numerical value consistent with the EFT estimate. We obtain the posterior distribution using a probabilistic integration method called Nested Sampling [Skilling06-BA] implemented in Python [2009MNRAS.398.1601F] that calculates both the posterior and the evidence.
The experimental data and EFT⋆ fits at LO, NLO, and NNLO are shown in Fig. 8. and its derivatives at threshold are given in Table 6. The quality of the fits, the consistency of the numerical values of the fit parameters with EFT assumptions, and the order-by-order improvement in the numerical value of the fitted capture cross section and factor give us confidence in the estimated theoretical uncertainty. We find the posterior probability distribution functions to be symmetric and well described by Gaussians, and as such we report the mean and standard deviation in Table 6. Table LABEL:table:normsBayesian gives the scaling factors for the experimental data sets determined from the NNLO fit along with their common mode errors.

### v2

@@ IX.1 Introduction
Radiative proton capture on (=3/2-) at solar energies proceeds through non-resonant capture to the ground state of (=2+). This capture occurs predominantly at separations well beyond the range of the strong interaction via transitions from and partial waves, as the amplitude and contributions from higher partial waves are negligible in the energy range of interest [1961NucPh..24...89C]. The importance of this reaction for the determination of the high energy solar neutrino spectrum and the experimental data reviewed by 1998RvMP...70.1265A inspired several experiments that, taken together, provided a consistent picture of the energy dependence of the reaction cross section, despite discrepancies in absolute scale. Prior to the review of 2011RvMP...83..195A, most direct measurements were performed using a radioactive target and an intense proton beam. In spite of the great care taken in evaluating the systematic energy-independent uncertainties, in the following referred to as common-mode errors (CMEs), a remarkable discrepancy persisted even in the most recent experiments, that was finally handled in 2011RvMP...83..195A by inflating the stated experimental uncertainties. Indirect measurements of the cross section of B based on Coulomb dissociation were considered, but not included in the final recommendation of 2011RvMP...83..195A for the zero-energy astrophysical S-factor . However, despite the experimental issues, the uncertainty on the recommended value of in 2011RvMP...83..195A is dominated by the theoretical contribution. Therefore, our analysis focused on these aspects (discrepancy of experimental data, indirect measurements, theoretical uncertainties), as discussed in the next sections.
@@ IX.2.1 Direct Measurements
The thorough review of published work done in 2011RvMP...83..195A investigated the influence of beam-target overlap, target stoichiometry, beam energy loss, and the backscattering of 8B recoils on CMEs and led to the selection of a homogeneous group of well-documented data sets. Since then, no further information became available on these aspects of direct measurements. The data that formed the basis of the 2011RvMP...83..195A recommendation were those of 2003PhRvC..67f5805B, 2003PhRvL..90b2501B, 1998PhRvL..80..928H, 2001PhRvL..86.3985H, 1999PhLB..462..237H, 2001NuPhA.696..219S, and the BE3 data set of 2003PhRvC..68f5803J. We considered the complete set of radiative capture measurements that were the basis of the 2011RvMP...83..195A recommendation for , where we adopted the revised BE3 data presented by 2010PhRvC..81a2801J, which were reanalyzed using a more sophisticated model of the target and more accurate stopping powers. In addition, we included in our analysis the only new experiment reported thereafter of 2022PhLB..82436819B, which used a radioactive ion beam and the recoil mass separator ERNA to detect the 8B recoils. While these new results contribute to the determination of with different, well controlled systematics [2018EPJA...54...92B], their large statistical uncertainty relative to the other experiments, due to the low counting statistics, limits their impact.
New scattering data have been published. Elastic and inelastic scattering cross sections at relative kinetic energies between 474 keV and 2.74 MeV were measured and wave scattering lengths were inferred by 2019PhRvC..99d5807P. At higher energies between 1.6 and 3.4 MeV, thick target elastic and inelastic scattering excitation function measurements were used to infer the existence of new resonances above 1.8 MeV [2013PhRvC..87e4617M]. These high lying resonances have no direct influence on because of their small partial widths in the channel and/or because they cannot be populated by s- or d-wave capture at low energy.
@@ IX.2.2 Indirect Measurements
In 2011RvMP...83..195A, Coulomb dissociation measurements based on the formalism developed by 1986NuPhA.458..188B were considered [PhysRevLett.83.2910, PhysRevLett.73.2680, KIKUCHI1997261, 1998EPJA....3..213K, PhysRevLett.86.2750, PhysRevC.63.065806, PhysRevLett.90.232501, 2006PhRvC..73a5806S], where in some cases the most recent publications supersede the previously published ones. Finally, the results of these experiments were not included in the determination of due to disagreements over whether the analyses of the various measurements properly accounted for the contributions of transitions. We disagree with these reservations and find the agreement of these measurements with radiative capture measurements significant. However, in order to include these consistent data in formulating our recommendation for , we inflated the CMEs of the Coulomb breakup measurements to adequately account for the uncertainty in the E2 components assumed in each analysis.
The common mode errors are obtained from a linear sum of those given in the manuscripts and an additional common mode error due to the estimated size of the component in each measurement. This additional error is applicable in one direction only, downward for the 1998EPJA....3..213K and 2006PhRvC..73a5806S analyses of the RIKEN and GSI measurements, respectively, and upward for the 2003PhRvC..68d5802D analysis of the Michigan State University (MSU) measurement, according to whether the component was included or excluded. Though the 2003PhRvC..68d5802D analysis did include a common mode error contribution of 2.5% due to the component (which represented 5% of the total cross section), we have included an additional 2.5% contribution to the MSU experiment. The size of the additional common mode errors for the RIKEN and GSI measurements was based on scaling this 2.5% according to the predicted size of the contribution to the cross section in these measurements relative to that in the MSU measurement, evaluated using first-order perturbation theory and the and S-factors calculated in a potential model of 8B [1996ZPhyA.356..293B]. Thus the additional single-sided common mode errors added to the GSI, MSU, and RIKEN measurements are -2.1%, 2.5%, and -4.3%, respectively.
@@ IX.3 Theory
The size of the theoretical contribution to the uncertainty in the 2011RvMP...83..195A estimate of was twice that of the experimental contribution. Since then, significant progress has been made. A new potential model calculation was performed by 2019NuPhA.983..175D. The large-scale computational demands of calculations using high-quality nucleon-nucleon (NN) and three-nucleon (3N) interactions beyond =5 had hindered the ab initio approach in the past. Nevertheless progress has been rapid and it has been possible to extend ab initio calculations to B, albeit without 3N forces. 2011PhLB..704..379N performed such a calculation using the no-core shell model in the continuum with a nucleon-nucleon interaction derived from chiral EFT at fourth order. The calculation converged with respect to the model space size and reproduces experimental data for the S-factor despite having virtually no free parameters (though the similarity renormalization group evolution parameter was tuned to reproduce the binding energy of 8B). The final value of =19.4(7) eV b agrees with the 2011RvMP...83..195A recommendation and, as will be shown later, with our analysis. More recently, a new approach was introduced to combine experimental measurements with ab initio predictions, resulting in an “ab initio-informed evaluation” Kravvaris23-PLB that arrived at a value of eV b.
@@
2015PhLB..751..535Z and 2018PhRvC..98c4616Z include such an excited contribution. 2022PhRvC.106a4601H provides a calculation without the contribution below and one with it above that also includes an contribution to the resonance. The two groups’ halo EFT calculations with the contributions differ in several respects that are elaborated in 2022PhRvC.106a4601H. Differences in the final bound state calculation affect the treatment of the short-range interactions and divergences in the EFTs. The interpretation of the asymptotic normalization constants (ANCs) in terms of elastic scattering parameters is affected. However, the momentum dependence of the capture cross section is not impacted by the ANCs, which are fitted to capture data. Thus there is no effect on the cross section. In the incoming channel, 2022PhRvC.106a4601H use a more general form of the -wave short-ranged interaction involving the core based on the low-energy symmetry that is included only at NNLO. In contrast, 2015PhLB..751..535Z and 2018PhRvC..98c4616Z include a short-ranged interaction at LO that could make a difference in the momentum dependence of the cross section. However, as shown in an order-by-order calculation [2022PhRvC.106a4601H], the short-ranged interaction is a sub-leading effect so the two halo EFTs should have similar low-momentum dependence. The third important difference is in the treatment of two-body currents, which affects EFT error estimates. 2015PhLB..751..535Z and 2018PhRvC..98c4616Z include this at NLO whereas 2022PhRvC.106a4601H estimate it to be an N3LO effect. Bayesian estimates of the two-body current [2015PhLB..751..535Z, 2018PhRvC..98c4616Z] are not strongly constrained by capture data, consistent with this being a higher-order contribution.
Another widely used approach for describing the cross sections of low-energy nuclear reactions is R-matrix theory. In R-matrix theory the 7Be(B cross section can be parametrized in terms of a few parameters representing either “real” or “background” poles. The real poles correspond to the observed states and resonances, while the background terms correspond to the “mean-field” effects that are of the non-resonant type. In the present case for example, a background pole in the 1- channel is needed so as to reproduce the non-resonant part of the capture. The parameters of the real pole corresponding to the bound state are determined by the binding energy and the two ANCs, while the 1+ resonance parameters are determined by its energy and partial decay widths (both proton and radiative). The R-matrix analysis of this reaction carried out by 1995NuPhA.588..693B arrived at a value of eV b; however, it was based on data that were excluded in 2011RvMP...83..195A and here, with the exception of those of 1983PhRvC..28.2222F.
@@ IX.4 Fitting Procedures
In 2011RvMP...83..195A, it was noted that the different data sets had similar energy dependences. The discrepancies among the data sets were primarily due to different absolute normalizations. We address the normalization issue by introducing scaling factors associated with the CMEs. This contrasts with the 2011RvMP...83..195A approach, in which the errors were inflated by the factor . The procedure for handling discrepant data sets adopted here is mathematically rigorous and aligns naturally with the method of estimating theoretical uncertainty discussed below. We prefer this approach to the inflation factor method used in 2011RvMP...83..195A, since the absolute scales of the different data sets can be used as constraints for each other, while there is no information available for a reevaluation of the CMEs. The procedure we used, described below, can be applied to both frequentist and Bayesian analyses.
Although our final recommendation is based on Bayesian analysis alone, both minimization (a frequentist approach) and Bayesian posterior probability distribution function evaluation were performed to fit EFTs to the data; in the R-matrix analysis, only the former was used. The likelihood function, defined as the conditional probability for the data given theoretical parameters , enters both fitting procedures. It can be derived from the expression given in 1994NIMPA.346..306D
@@ IX.5 Data Analysis and Determination
Bayesian fits were performed with halo EFT expressions from both 2020JPhG...47e4002Z and 2018EPJA...54...89H that are compatible with the various fits. The Bayesian analysis had advantages over the fits in exploring parameter space and in determining realistic uncertainties [PhysRevLett.122.232502].
The capture proceeds through the and channels. The reaction is known to be peripheral, resulting in a that has a sub-leading dependence on the strong interaction in the initial state [2000PhRvC..62f5803B]. This behavior was shown to persist away from the threshold in halo EFT, in which for capture the strong interaction in the incoming -wave channel only contributes at NNLO.
@@
The Bayesian fits involved six theoretical parameters and 11 normalization constants for the data sets. The priors for the normalization constants were as described earlier. The priors for the theoretical parameters are based on the underlying EFT assumptions [2022PhRvC.106a4601H] about the expansion in a small ratio where represents the low momentum of infrared physics and the cutoff of the theory. The and scattering lengths fm and fm [2019PhRvC..99d5807P] are taken as inputs. For capture at LO, the only fit parameter, the effective momentum is assumed to scale as and accordingly we assume a uniform prior for between -100 MeV/ and 1.5 MeV/, where the upper limit is set by the physical constraint that the ANC squares have to be positive. At NLO, the ANC2 is a fit parameter whose prior is taken from a normal distribution determined by experiment. At NNLO, we use the ab initio calculation of [2018PhRvC..98c4616Z] to draw the ANC2 from a normal distribution. A ratio of the -wave scattering lengths in the coupled-channel - that is assumed to be in the EFT is drawn from a normal distribution with a mean of 0 and a standard deviation of 10 to cover a wide range. The capture is dominated by the transition and it requires three parameters: a -wave scattering volume , an effective momentum and a two-body current to regulate divergences. 2022PhRvC.106a4601H determined and from the the narrow resonance energy and width. Here, we kept fixed while drawing from a uniform distribution between -150 and -50 MeV. The Bayesian fit gave . The two-body coupling was drawn from a uniform distribution and the fits gave a numerical value consistent with the EFT estimate. We obtain the posterior distribution using a probabilistic integration method called Nested Sampling [Skilling06-BA] implemented in Python [2009MNRAS.398.1601F] that calculates both the posterior and the evidence.
The experimental data and EFT⋆ fits at LO, NLO, and NNLO are shown in Fig. 8. and its derivatives at threshold are given in Table 7. The quality of the fits, the consistency of the numerical values of the fit parameters with EFT assumptions, and the order-by-order improvement in the numerical value of the fitted capture cross section and S-factor give us confidence in the estimated theoretical uncertainty. We find the posterior probability distribution functions to be symmetric and well described by Gaussians, and as such we report the mean and standard deviation in Table 7. Table LABEL:table:normsBayesian gives the scaling factors for the experimental data sets determined from the NNLO fit along with their common mode errors.